How do I find a quadratic equation that has -6 and 4 as solutions in the form of ax^2+bx+c=0
Answer by ankor@dixie-net.com(22740)

How do I find a quadratic equation that has -6 and 4 as solutions in the form of ax^2+bx+c=0
:
Working back
x = -6, x = 4
(x+6)(x-4) = 0
FOIL
x^2 - 4x + 6x - 24 = 0
:
y = x^2 + 2x - 24